FINDING SAMPLE SIZE.
Q1. margin of error:$125, confidence level:95%,standard deviation=$500

Answer by stanbon(48568)

FINDING SAMPLE SIZE.
Q1. margin of error:$125, confidence level:95%,standard deviation=$500
-----------
E=z*(sigma/sqrt(n))
sqrt(n) = z*(sigma)/E
sqrt(n) = 1.96(500)/125
sqrt(n) = 7.84
n=61 (rounded to whole number)
========
Cheers,
Stan H.